# Neural Wave Functions for Superfluids

### v3

@@ IV.2 AGPs FermiNet with Unpaired States
$=\sum_{k}\omega_{k}\det\left(\begin{array}[]{cccc}\varphi^{k}(\mathbf{r}^{\uparrow}_{1},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{1},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{1};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}\})\\ \varphi^{k}(\mathbf{r}^{\uparrow}_{2},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{2},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{2};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}\})\\ \vdots&\ddots&\vdots&\vdots\\ \varphi^{k}(\mathbf{r}^{\uparrow}_{p+1},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{p+1},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{p+1};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}\})\end{array}\right)$ (43)
@@ V Results
The power of the FermiNet AGPs Ansatz may be demonstrated by studying the UFG. The Hamiltonian is
@@
is the modified Pöschl-Teller potential, which is widely used in variational and diffusion QMC simulations [16, 17, 18, 19, 20, 21, 22, 23]. It would be better to use a delta function interaction with an infinite $s$-wave scattering length, but it is difficult to simulate systems with delta-like potentials using QMC methods. Thus, a finite but short-ranged interaction is typically used. The $s$-wave scattering length of the Pöschl-Teller potential diverges when $v_{0}=1$. By changing the value of $\mu$ at fixed $v_{0}=1$, it is possible to vary the effective range of the interaction, $r_{e}=2/\mu$, whilst holding the $s$-wave scattering length infinite.
@@
We choose to study a system with density parameter $r_{s}=1$, where $r_{s}$, the radius of a sphere that contains one particle on average, provides a convenient measure of the inter-particle distance. Throughout this work, we employ the dimensionless system based on Hartree atomic units: the unit of length is the Bohr radius, $a_{0}$, and the unit of energy is the Hartree. To ensure that the range of the interaction is small compared with the inter-particle separation, we set $\mu=12$ ($r_{e}=1/6$), keeping $v_{0}=1$ to ensure that the scattering length remains infinite 444 Although the range of the potential is finite, which introduces a second length scale even at unitarity, most of the results in the literature have been obtained with the same value of $\mu$. This allows us to compare our results with theirs. To obtain exact results, one has to study the limit as $\mu\rightarrow\infty$, but converging the simulations becomes more difficult as $\mu$ increases. . We have also simulated the system with $k_{F}=1$ (equivalent to $r_{s}=(9\pi/4)^{1/3}\approx 1.92$) and $\mu=12$ to compare with the fixed-node diffusion Monte Carlo (FN-DMC) result from Forbes et al. [21].
@@
We use both the original FermiNet and the AGPs FermiNet to study the unitary fermi gas from $N=4$ to $N=38$ particles in a cubic box subject to periodic boundary conditions. The same network size, number of determinants, and number of training iterations are used for both Ansätze. The FermiNet orbitals are given by Eq. (9) without the bias term. The FermiNet geminal used for systems containing from $4$ to $28$ particles is the one defined as in Eq. (17). Including contributions from the two-electron stream improves the optimization rate and can achieve a slightly lower variational energy in larger systems, so Eq. (19) was used for systems of $29$ to $38$ particles. The inclusion of plane-wave envelopes as defined in Eq. (12) also improves the optimization rate.
@@
A comparison of the ground-state energy expectation values given by the two Ansätze is shown in Fig. 0(a). The conventional FermiNet Ansatz, which consists of a linear combination of block-diagonal determinants of FermiNet orbitals, performs well when the number of particles $N$ is smaller than around $10$, but the AGPs FermiNet is much superior in larger systems. It is clear that the original FermiNet Ansatz has difficulties learning the ground states of large paired systems.
@@
In systems containing an odd number of particles, one must be left unpaired. This raises the energy a little and explains the zigzag shape of Fig. 0(a). The odd-even staggering is lost for larger systems with the original FermiNet Ansatz, indicating the absence of pair formation [16, 56]. The original FermiNet fails to learn the superfluid state. For the AGPs FermiNet, by contrast, the amplitude of the odd-even zigzag remains constant, superposed on the linear increase with $N$ expected of any extensive quantity.
@@
Another comparison between the two Ansätze is shown in Fig. 0(b) depicting the ratio of the interacting and non-interacting energies per particle, known as the Bertsch parameter [4] and defined in Eq. (1), as a function of $N$. All FermiNet energies are variational and the non-interacting energies are exact, so the AGPs FermiNet, for which the Bertsch parameter is lower by up to around 30%, is the much better of the two Ansätze.
@@
(a) The total energy of the UFG simulation cell, measured in units of the free Fermi gas energy $E_{\text{FG}}$. The original FermiNet Ansatz begins to fail when $N\gtrapprox 10$.
@@
Figure 1: Comparison between results obtained using the AGPs FermiNet and the original FermiNet for different numbers of particles, $N$, with $r_{s}=1$ and $\mu=12$. All simulations used 32 determinants, 300,000 optimization steps, and the same hyperparameters, which are detailed in the Appendix.
We next compare our results with state-of-the-art FN-DMC results from Forbes et al. [21] shown in Fig. 2 for the case $k_{F}=1$ and $\mu=12$.
Figure 2: Comparison of the system-size dependent values of the Bertsch parameter, $\xi$, as calculated using the AGPs FermiNet and FN-DMC, with $k_{F}=1$ and $\mu=12$. According to the variational principle, lower values are better. The error bars on the AGPs FermiNet results are smaller than the sizes of the crosses. Inset: difference between the AGPs and FN-DMC values of the Bertsch parameter.
The AGPs FermiNet achieves a lower energy per particle than FN-DMC for all system sizes except for $N=4$ and $N=6$. The dependence of the Bertsch parameter on system size is also smoother when calculated with the AGPs FermiNet.
@@
Here $a$ is the scattering length of the interaction, which is infinite in the UFG. In this limit, $\Delta_{\text{BCS}}=1.804E_{\text{FG}}$ and $\Delta_{\text{Gorkov}}=0.815E_{\text{FG}}$, where $E_{\text{FG}}=\frac{3}{5}\frac{\hbar^{2}k_{F}^{2}}{2m}$ is the average energy per particle of an unpolarized non-interacting Fermi gas and $e$ is Euler’s number. The UFG is a strongly coupled system, so the BCS and Gorkov estimates of the gap need not be accurate.
The striking collapse of the pairing gap with increasing system size shows that the original FermiNet Ansatz struggles to describe paired states in systems of more than $10$ particles. The AGPs FermiNet Ansatz behaves much better, although the oscillations with system size suggest that significant finite-size errors remain even for the largest systems simulated.
Figure 3: Pairing gaps calculated with the original FermiNet and the AGPs FermiNet for different numbers of particles $N$, with $r_{s}=1$ and $\mu=12$.
@@
$\rho_{\uparrow\downarrow}^{(2)}(\mathbf{r}_{1},\mathbf{r}_{2};\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=\langle\hat{\psi}^{\dagger}_{\uparrow}(\mathbf{r}_{1})\hat{\psi}^{\dagger}_{\downarrow}(\mathbf{r}_{2})\hat{\psi}_{\downarrow}(\mathbf{r}^{\prime}_{2})\hat{\psi}_{\uparrow}(\mathbf{r}_{1}^{\prime})\rangle,$ (49)
@@
The one-body density matrix (OBDM), by contrast, tends to zero in the $r\rightarrow\infty$ limit [58]. A full discussion of the methods used to evaluate the condensate fraction in QMC simulations can be found in the Appendix and the CASINO manual [59].
After fully training both the original FermiNet and the AGPs FermiNet for the $N=38$ particle system, we used the resulting neural wave functions to compute the quantity $\frac{\Omega^{2}}{N_{\uparrow}}\rho_{\uparrow\downarrow}^{(2)TR}(r)$. The results are shown in Fig. 4, which provide further evidence that the original FermiNet fails to converge to the superfluid ground state; the quantity $\frac{\Omega^{2}}{N_{\uparrow}}\rho_{\uparrow\downarrow}^{(2)TR}(r)$ appears to be approaching zero in the large pair-separation limit, implying that the condensate fraction is also zero. The same quantity for the AGPs FermiNet approaches a finite value which we esitmated to be roughly $c=0.44(1)$ using the eight data points with separations $r/r_{s}\geq 2.0$. This value is consistent with previous estimations from experiments and the most recent AFMC value from [22]. Some of the results are summerized in Table (1)
@@
Our estimation 0.44(1)
FN-DMC [60] 0.57(2)
FN-DMC and VMC with extrapolation [18] 0.51
FN-DMC with effective range extrapolation [61] 0.56(1)
AFMC [22] 0.43(2)
@@
Table 1: The estimations of the superfluid condensate fraction at unitarity using various methods.
Figure 4: Comparing the TBDM estimators between the AGPs FermiNet and the original FermiNet with $N=38$, $r_{s}=1$ and $\mu=12$.
An important advantage of VMC methods is that almost any expectation value, including any reduced density matrix, may be estimated without bias. The same is not true of FN-DMC simulations, which sample the wave function instead of its square modulus and produce biased “one-sided” estimates of the expectation values of operators that do not commute with the Hamiltonian [14]. Thus, there are very few unbiased and accurate first-principles calculations of the condensate fraction. Our approach provides a solution to this problem and a more accurate way to estimate general expectation values.

### v4

@@ IV.2 AGPs FermiNet with Unpaired States
$=\sum_{k}\omega_{k}\det\left(\begin{array}[]{cccc}\varphi^{k}(\mathbf{r}^{\uparrow}_{1},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{1},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{1};\{\mathbf{r}^{\uparrow}_{/1}\};\{\mathbf{r}^{\downarrow}\})\\ \varphi^{k}(\mathbf{r}^{\uparrow}_{2},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{2},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{2};\{\mathbf{r}^{\uparrow}_{/2}\};\{\mathbf{r}^{\downarrow}\})\\ \vdots&\ddots&\vdots&\vdots\\ \varphi^{k}(\mathbf{r}^{\uparrow}_{p+1},\mathbf{r}^{\downarrow}_{1};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}_{/1}\})&\cdots&\varphi^{k}(\mathbf{r}^{\uparrow}_{p+1},\mathbf{r}^{\downarrow}_{p};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}_{/p}\})&\phi^{k\uparrow}_{1}(\mathbf{r}^{\uparrow}_{p+1};\{\mathbf{r}^{\uparrow}_{/p+1}\};\{\mathbf{r}^{\downarrow}\})\end{array}\right).$ (35)
where $\varphi^{k}(\mathbf{r}^{\uparrow}_{i},\mathbf{r}^{\downarrow}_{j};\{\mathbf{r}^{\uparrow}_{/i}\};\{\mathbf{r}^{\downarrow}_{/j}\})$ can either be defined as Eq. (17) or Eq. (19), and $\phi^{k\uparrow}_{i}(\mathbf{r}^{\uparrow}_{j};\{\mathbf{r}^{\uparrow}_{/j}\};\{\mathbf{r}^{\downarrow}\})$ is defined in Eq. (9). In practice, we generate the required number of FermiNet geminal and orbital parameters in batch at the orbital projection layer. For example, one FermiNet geminal and one FermiNet orbital are generated per determinant for a system with one extra spin-up particle.
@@ V Results
The power of the AGPs FermiNet Ansatz may be demonstrated by studying the UFG. The Hamiltonian is
@@
is the modified Pöschl-Teller potential, which is widely used in variational and diffusion QMC simulations [16, 17, 18, 19, 20, 21, 22, 23]. It would be preferable to use a delta function interaction with an infinite $s$-wave scattering length, but it is difficult to simulate systems with delta-like potentials using QMC methods. Thus, a finite but short-ranged interaction is typically used. The $s$-wave scattering length of the Pöschl-Teller potential diverges when $v_{0}=1$. By changing the value of $\mu$ at fixed $v_{0}=1$, it is possible to vary the effective range of the interaction, $r_{e}=2/\mu$, whilst holding the $s$-wave scattering length infinite.
@@
We choose to study a system with density parameter $r_{s}=1$, where $r_{s}$, the radius of a sphere that contains one particle on average, provides a convenient measure of the inter-particle distance. Throughout this work, we employ the dimensionless system based on Hartree atomic units: the unit of length is the Bohr radius, $a_{0}$, and the unit of energy is the Hartree. To ensure that the range of the interaction is small compared with the inter-particle separation, we set $\mu=12$ ($r_{e}=1/6$), keeping $v_{0}=1$ to ensure that the scattering length remains infinite 555 Although the range of the potential is finite, which introduces a second length scale even at unitarity, most of the results in the literature have been obtained with the same value of $\mu$. This allows us to compare our results with theirs. To obtain exact results, one has to study the limit as $\mu\rightarrow\infty$, but converging the simulations becomes more difficult as $\mu$ increases. . We have also simulated the system with $k_{F}=1$ (equivalent to $r_{s}=(9\pi/4)^{1/3}\approx 1.92$) and $\mu=12$ to compare with the fixed-node diffusion Monte Carlo (FN-DMC) result from Forbes et al. [21].
@@
We use both the Slater FermiNet with multiple block-diagonal determinants and the AGPs FermiNet with multiple geminals to study the unitary Fermi gas from $N=4$ to $N=38$ particles in a cubic box subject to periodic boundary conditions, as well as AGPs FermiNet on the $N=66$ system 666 As previous works [22, 74, 75, 18] have shown, the ground-state energy of the $N=66$ particle UFG is the closest to the ground-state energy in the thermodynamic limit even among larger systems. This can also be seen by comparing the ratio of $E_{FG}$ and the energy per particle from summing over the energies occupied by the particles in the discretized $k$-space, where this ratio is just $0.5\%$ away from unity for $N=66$. . The same network size, number of determinants, and number of training iterations are used for both Ansätze. The FermiNet orbitals are given by Eq. (9) without the bias term. The FermiNet geminal used for systems containing from $4$ to $28$ particles is the one defined in Eq. (17). Unless otherwise stated, all calculations used a linear combination of 32 determinants or 32 geminals. Including contributions from the two-electron stream improves the optimization rate and can achieve a slightly lower variational energy in larger systems, so Eq. (19) was used for systems of $N\geq 29$. The inclusion of plane-wave envelopes as defined in Eq. (12) also improves the optimization rate. For molecular and electron gas systems, we have found that the bias term in the FermiNet orbital projection (Eq. 9) does not affect the accuracy or optimization of the model. We hence set the bias term, $g^{k\alpha}_{i}$ and $g^{k}$ as appropriate, to zero for all calculations presented here.
@@
A comparison of the ground-state energy expectation values given by the two Ansätze is shown in Fig. 1(a). The Slater FermiNet, which consists of a linear combination of block-diagonal determinants of FermiNet orbitals, performs well when the number of particles $N$ is smaller than around $10$, but the AGPs FermiNet is much superior in larger systems. It is clear that the Slater FermiNet Ansatz has difficulties learning the ground states of large paired systems 777 All calculations presented in this section are performed using the Slater FermiNet with block-diagonal determinants. A small set of calculations are performed with the Slater FermiNet with dense determinants, which are presented in the Appendix section (Appendix D). In short, we did found slight improvements over the block-diagonal Slater FermiNet results, but it is still qualitatively wrong and is also much worse than the AGPs FermiNet results. .
@@
In systems containing an odd number of particles, one must be left unpaired. This raises the energy a little and explains the zigzag shape of Fig. 1(a). The odd-even staggering is lost for larger systems with the Slater FermiNet Ansatz, indicating the absence of pair formation [16, 59]. The Slater FermiNet fails to learn the superfluid state. For the AGPs FermiNet, by contrast, the amplitude of the odd-even zigzag remains constant, superposed on the linear increase with $N$ expected of any extensive quantity.
@@
Another comparison between the two Ansätze is shown in Fig. 1(b), which depicts the ratio of the interacting and non-interacting energies per particle, known as the Bertsch parameter [4] and defined in Eq. (1), as a function of $N$. All FermiNet energies are variational and the non-interacting energies are exact, so the AGPs FermiNet, for which the Bertsch parameter is lower by up to around 30%, is much the better of the two Ansätze.
@@
(a) The total energy of the UFG simulation cell, measured in units of the free Fermi gas energy $E_{\text{FG}}$. The Slater FermiNet Ansatz begins to fail when $N\gtrapprox 10$.
@@
Figure 1: Comparison between results obtained using the AGPs FermiNet and the Slater FermiNet for different numbers of particles, $N$, with $r_{s}=1$ and $\mu=12$. All simulations used 32 determinants, 300,000 optimization steps, and the same hyperparameters, which are detailed in the Appendix.
We next compare our results with the state-of-the-art FN-DMC results of Forbes et al. [21], shown in Fig. 2 for the case $k_{F}=1$ and $\mu=12$.
Figure 2: Comparison of the system-size dependent values of the Bertsch parameter, $\xi$, as calculated using the AGPs FermiNet and FN-DMC, with $k_{F}=1$ and $\mu=12$. According to the variational principle, lower values are better. The error bars on the AGPs FermiNet results are smaller than the sizes of the crosses. Inset: difference between the AGPs and FN-DMC values of the Bertsch parameter. The errors in the inset are obtained by adding the standard errors of the FN-DMC and AGPs FermiNet results in quadrature. The latter are obtained by computing the standard error of the MCMC-averaged Bertsch parameter accumulated over 50,000 inference steps.
The AGPs FermiNet achieves a lower energy per particle than FN-DMC for all system sizes except for $N=4$ and $N=6$. The dependence of the Bertsch parameter on system size is also smoother when calculated with the AGPs FermiNet 888 Limited by the scope of this work, we did not proceed to carry out a similar set of calculations done by Forbes et al. [21] for extrapolating the results to $\mu\rightarrow\infty$. . A full training curve of $N=66$ with comparison to the FN-DMC energy can be found in Appendix F.
@@
Here $a$ is the scattering length of the interaction, which is infinite in the UFG. In this limit, $\Delta_{\text{BCS}}=1.804E_{\text{FG}}$ and $\Delta_{\text{Gorkov}}=0.815E_{\text{FG}}$, where $E_{\text{FG}}=\frac{3}{5}\frac{\hbar^{2}k_{F}^{2}}{2m}$ is the average energy per particle of an unpolarized non-interacting Fermi gas and $e$ is Euler’s number 999 Although writing the pairing gap in units of the Fermi energy would be more suitable as it is a one-body quantity, we choose to express it in units of $E_{FG}$ instead, because i) we want to be consistent with the units used in the paper by J. Carlson et al. [16], which also used $E_{FG}$ as the energy unit for the pairing gap, and ii) to keep the units consistent with other quantities in this work. . The UFG is a strongly coupled system, so the BCS and Gorkov estimates of the gap need not be accurate.
The striking collapse of the pairing gap with increasing system size shows that the Slater FermiNet Ansatz struggles to describe paired states in systems of more than $10$ particles. The AGPs FermiNet Ansatz behaves much better, although the oscillations with system size suggest that significant finite-size errors remain even for the largest systems simulated.
Figure 3: Pairing gaps calculated with the Slater FermiNet and the AGPs FermiNet for different numbers of particles $N$, with $r_{s}=1$ and $\mu=12$.
@@
$\rho_{\uparrow\downarrow}^{(2)}(\mathbf{r}_{1},\mathbf{r}_{2};\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=\langle\hat{\psi}^{\dagger}_{\uparrow}(\mathbf{r}_{1})\hat{\psi}^{\dagger}_{\downarrow}(\mathbf{r}_{2})\hat{\psi}^{\phantom{\dagger}}_{\downarrow}(\mathbf{r}^{\prime}_{2})\hat{\psi}^{\phantom{\dagger}}_{\uparrow}(\mathbf{r}_{1}^{\prime})\rangle,$ (41)
@@
The one-body density matrix, by contrast, tends to zero in the $r\rightarrow\infty$ limit [63]. A full discussion of the methods used to evaluate the condensate fraction in QMC simulations can be found in the Appendix and the CASINO manual [64].
After fully training both the Slater FermiNet and the AGPs FermiNet for the $N=38$ particle system, we used the resulting neural wave functions to compute the quantity $\frac{\Omega^{2}}{N_{\uparrow}}\rho_{\uparrow\downarrow}^{(2)TR}(r)$. The results are shown in Fig. 4, which provide further evidence that the Slater FermiNet fails to converge to the superfluid ground state; the quantity $\frac{\Omega^{2}}{N_{\uparrow}}\rho_{\uparrow\downarrow}^{(2)TR}(r)$ appears to be approaching zero in the large pair-separation limit, implying that the condensate fraction is also zero. The same quantity for the AGPs FermiNet approaches a finite value which we estimated to be roughly $c=0.44(1)$ using the eight data points with separations $r/r_{s}\geq 2.0$. This value is consistent with previous estimations from experiments and the most recent AFMC value from [22] (Table 1).
In addition, we also estimated the condensate fractions for the $N=66$ UFG at a fixed $\mu=12$ with two different densities: $r_{s}=1$ ($k_{F}r_{e}=0.32$) and $k_{F}=1$ ($k_{F}r_{e}=0.17$), respectively. We compute the quantity $\frac{\Omega^{2}}{N_{\uparrow}}\rho_{\uparrow\downarrow}^{(2)TR}(r)$ at five sequentially-spaced separations $r$ near $r=L/2$, where the quantity has approached its asymptotic value. We then take the average of the five data points to get estimated values of the condensate fraction. 101010 Due to limited resources, we only compute five data points at far separations instead of the whole range of separations from $r=0$ to $r=L/2$ for $N=66$. . Our estimate of the condensate fraction for $N=66$ is $c=0.42(1)$ at $k_{F}r_{e}=0.32$, and $c=0.52(1)$ at $k_{F}r_{e}=0.17$, which are both consistent with the experiments. The results are summarized in Table 1.
@@
Our estimate for $N=38$ at $k_{F}r_{e}=0.32$ 0.44(1)
Our estimate for $N=66$ at $k_{F}r_{e}=0.32$ 0.42(1)
Our estimate for $N=66$ at $k_{F}r_{e}=0.17$ 0.52(1)
FN-DMC for $N=38$ at $k_{F}r_{e}=0.03$ [66] 0.61(2)
FN-DMC for $N=66$ at $k_{F}r_{e}=0.03$ [66] 0.57(2)
FN-DMC for $N=128$ with VMC extrapolation at $k_{F}r_{e}=0.32$ [18] 0.51(1)
FN-DMC with $k_{F}r_{e}\rightarrow 0$ extrapolation for $N=66$ [67] 0.56(1)
AFMC with $k_{F}r_{e}\rightarrow 0$ extrapolation for $N=66$ [22] 0.43(2)
@@
Table 1: Estimates of the superfluid condensate fraction at unitarity using various methods. The quantity $k_{F}r_{e}$ is a dimensionless number, indicating the deviation of the simulated system from a perfect UFG with zero-range interaction.
Figure 4: Comparison of the TBDM estimators calculated using the AGPs FermiNet and the Slater FermiNet with $N=38$, $r_{s}=1$ and $\mu=12$. The error bars show the standard error of the TBDM estimator, accumulated over 2,000 inference steps. Most of the error bars are so small that they are obscured the symbols.
Although VMC methods are generally considered to be less accurate than FN-DMC methods, an important advantage of VMC methods is that almost any expectation value, including any reduced density matrix, may be estimated without bias. The same is not true of FN-DMC simulations, which sample the wave function instead of its square modulus and produce biased “one-sided” estimates of the expectation values of operators that do not commute with the Hamiltonian [14]. Thus, there are very few unbiased and accurate first-principles calculations of the condensate fraction. Our approach, having both the advantages of VMC and surpassing the accuracy of DMC, provides solutions to these problems and a more accurate way to estimate general expectation values.
Finally, we study how the number of block-diagonal determinants required to achieve a given accuracy scales with the number of particles in the system. We choose six even-particle systems from $N=4$ to $N=14$ and compare the energies obtained using linear combinations of multiple block-diagonal FermiNet Slater determinants against energies obtained using a single-determinant (and thus single-geminal) FermiNet AGPs wave function. All other hyperparameters are as given in Table 2 (Appendix B.1).
Figure 5: Number of block-diagonal determinants required for the multi-determinant Slater FermiNet energy to fall within 5% and 10% of the energy obtained using a single-determinant/single-geminal AGPs FermiNet.
In Fig. 5, we show that the number of block-diagonal determinants required to achieve a given percentage accuracy increases approximately exponentially with the number of particles. Plots for each individual system, along with a more detailed discussion, can be found in Appendix C. These results suggest that multi-determinant Slater FermiNet wave functions constructed using a neural network of fixed size are incapable of describing the ground state of the UFG accurately unless the number of block-diagonal determinants rises exponentially with system size. Hence, in practice, the AGPs FermiNet is required for studying paired systems.